# Beyond Pipelines: A Survey of the Paradigm Shift toward Model-Native Agentic AI

### v1

@@ 7.1.2. Reflection
This conflict is actually a common phenomenon in the early stage of a technological revolution. History offers a compelling parallel. The early days of the Internet saw it evolve from a basic infrastructure for information access into a ubiquitous utility, with its power encapsulated in websites and applications. As illustrated in the top of Fig. 11, its architectural evolution clearly demonstrates a dynamic shift in the balance between engineering and intelligence: from an early, purely engineering stage of hand-coded HTML/CSS to a dynamic interaction stage where applications like search engines began to incorporate statistics-based intelligence. This has finally culminated in the present day, where cloud services and MLOps have standardized the underlying engineering, allowing AI-driven personalization and recommendation to become the core value of web applications.
The evolution of web applications suggests that the current engineering-heavy agentic AI is transitory. As the underlying technologies discussed in this survey mature, the applications built upon them will likely follow a similar path. As illustrated in the bottom of Fig. 11, we envision the development of these agent systems in three stages:
@@
    During this period, scaffolding frameworks such as LangChain, AutoGen, and AgentCore will continue to play an important role. They will act as intermediate layers that bridge the gap between model-native intelligence and real-world deployment.
@@
Figure 11. Comparison of the evolution of web applications and agent systems
As agentic models become more self-sufficient, the role of the system layer will undergo a critical transformation: from compensating for the model’s capability deficits to providing foundational system support. Today, system engineering is primarily focused on patching the limitations of agentic models. In the future, the system layer will provide the essential, high-level infrastructure required for a robust and scalable agent ecosystem. The possible functions include identity management, resource allocation of compute and tools, standardized agent protocols, and granular permissions management to ensure safe and compliant operation.

### v2

@@ 7.1.2. Reflection
This conflict is actually a common phenomenon in the early stage of a technological revolution. History offers a compelling parallel. The architectural evolution of Internet clearly demonstrates a dynamic shift in the balance between engineering and intelligence: from an early, purely engineering stage of hand-coded HTML/CSS to a dynamic interaction stage where applications like search engines began to incorporate statistics-based intelligence. This has finally culminated in the present day, where cloud services and MLOps have standardized the underlying engineering, allowing AI-driven personalization and recommendation to become the core value of web applications.
The evolution of web applications suggests that the current engineering-heavy agentic AI is transitory. As the underlying technologies discussed in this survey mature, the applications built upon them will likely follow a similar path. We envision the development of these agent systems in three stages:
@@
    During this period, scaffolding frameworks such as LangChain, AutoGen, AgentCore and Agent Builder will continue to play an important role. They will act as intermediate layers that bridge the gap between model-native intelligence and real-world deployment.
@@
As agentic models become more self-sufficient, the role of the system layer is undergoing a critical transformation: from compensating for capability deficits to providing foundational ecosystem support. This transition is actually a recurring pattern which is evident in the development of LLM capabilities over the past few years, as illustrated in Fig. 11. At each stage, as a new capability becomes model-native, the system layer’s role evolves from “compensator” to “supporter”:
  • Dialogue: Traditional dialogue systems often split into modules like intent understanding, dialogue management, and response generation. With the rise of ChatGPT, the model layer internalized this entire flow. Consequently, the system layer’s role shifted from capability compensation to ecosystem support: providing external data access and context management, exemplified by frameworks like LangChain and LlamaIndex.
  • Planning : As models like o1 began to internalize reasoning through RL, the system layer’s role transformed to support this new training paradigm by providing robust Reinforcement Fine-Tuning (RFT), and evaluation frameworks, such as those from OpenAI and Statsig.
  • Tool Use : As models like o3 internalize the ability to plan and execute tool calls, the system layer’s role moves away from workflow orchestration and toward providing a standardized ecosystem for tool integration, such as MCP-based tool transformation and scalable server integration (e.g., Fixie.ai, Compositio).
Figure 11. The evolving role of the system layer (2022-2025): from capability compensation to ecosystem support
This established pattern provides a view for understanding the system layer’s future evolution, which will be driven by the continued internalization of other agentic capabilities, as illustrated in Fig. 12. The next stages of this evolution are likely to be:
  • Memory: The system layer will shift from its current role of compensating for the model’s statelessness with external vector databases and context engineering (e.g., Letta, Pinecone). It will transition to supporting the model’s native memory by providing a foundational ecosystem for cross-platform memory synchronization, data compliance, and privacy controls.
  • Multi-Agent Collaboration: The system layer will move beyond compensating for a lack of collaboration capability with pre-defined, handcrafted roles and externally orchestrated workflows (e.g., CrewAI, LangGraph). It will transform to support emergent collaboration by providing essential services like environment and resource scheduling, managing inter-agent communication, and offering a platform for collaborative evaluation.
Figure 12. The future role of the system layer (2025-): ecosystem support and foundational AgentOps
Ultimately, as core agentic capabilities are progressively internalized, the system layer will evolve from its current state as a collection of capability-specific patches into a foundational AgentOps infrastructure. This future system layer will focus on providing domain-agnostic, scalable services essential for a robust agent ecosystem, evolving existing infrastructure to meet new, autonomous demands:
  • Agent Identity Management: The paradigm for human/device identity and authorization (e.g., Okta) will be extended to manage agent identity, enabling secure, auditable authentication for autonomous entities.
  • Agent Communication: The infrastructure for human-to-human communication (e.g., Twilio) will evolve to support high-bandwidth, structured agent-to-agent communication protocols.
  • Agent Execution Observability: The focus will expand from traditional system-level observability (e.g., Datadog) to the much more complex challenge of agent execution process observability, which involves tracing and debugging the emergent, non-deterministic decision-making paths of autonomous agents.